# Hyperdimensional Probe: Decoding LLM Representations via Vector Symbolic Architectures

### v1

@@ 5.3 From input-completion tasks to question-answering
This dataset evaluates extractive question answering – i.e. each answer is a text span within the input context – through questions generated by crowdworkers over Wikipedia articles. It fits our concept-focused probing objective, as its questions/answers map to concepts presented in the given context. This elicits the language model to focus on those concepts, allowing us to benchmark the extracted concepts against features derived from both questions and answers. We extract input features based on lexical semantics, exploiting WordNet (Miller, 1995) and DBpedia (Lehmann et al., 2015) as knowledge bases. The input representations (Section˜4.2) are then created using bundling, such as:
“What was the name of the ship that Napoleon sent to the Black Sea? (5)
@@
We generate 693,886 training inputs $\mathcal{Q}$ by incrementally considering questions with their corresponding features (see also Appendix˜I). Our trained encoder (Section˜4.4) achieves a test-set cosine similarity of 0.44, and a binary accuracy of 0.70 with Llama3.1. For our experiments, we consider 10,000 sampled questions $\bar{\mathcal{Q}}$, each now prefixed with its contextual text. We then probe the model’s final state before and after the autoregressive text generation, and extract concepts encoded in VSA encodings (Section˜4.5) by comparing them directly with the codebook $\Phi$, as binding is not involved. LLM performance on $\bar{\mathcal{Q}}$ shows an average F1 score of 0.69 $\pm$ 0.38, exact match of 0.52 $\pm$ 0.5, with 68% of outputs mentioning the target answers. On average, our probe extracts three concepts both before and after the model’s text generation.
@@
Observing concept drift.
We evaluate semantic-based concept relevance by computing cosine similarity between concept embeddings and question-answer features. The average similarity of extracted concepts related to the question decreases after text generation: by 4.8% for the entire sample and by 8.0% in the LLM error subset (32% of the sample), while no significant differences are observed prior to generation (Figure˜5; left). For answer-related concepts, overall no change is observed before and after text generation, but the LLM error subset shows a slight increase (+3.2–3.5%; Figure˜5, right). This suggests that LLM failures may stem from losing focus on the question rather than from a lack of answer-related knowledge. This hypothesis is supported by a weak positive Spearman correlation $(0.2$ with a p-value of $1e^{-99}$; Appendix˜K) between LLM’s F1 score and the proportion of question-related concepts extracted after text generation. For example, for the SQuAD query “What do laboratories try to produce hydrogen from?” (target answer: “solar energy and water”), the model erroneously outputs “water and heat” (F1 = 0.57). Before the model’s text generation, our proposed approach extracts the concepts $\phi_{\textrm{try}}$, $\phi_{\textrm{produce}}$, $\phi_{\textrm{hydrogen}}$ (question) and $\phi_{\textrm{solar}}$, $\phi_{\textrm{water}}$ (answer); after generation, the question-related concept set reduced to $\phi_{\textrm{produce}}$ and the answer set gained the concept $\phi_{\textrm{energy}}$. While answer-related concepts refined, the model lost focus on the subject hydrogen, drifting toward a generic notion of production and ultimately an error.
@@ 6 Conclusions
This work offers empirical evidence supporting the hypothesis that LLM embeddings can be accurately represented using Vector Symbolic Architectures (VSAs), combining ideas from symbolic representations and neural probing. Our Hyperdimensional probe is found to effectively extract human-interpretable information from latent representations of LLMs, as reported in Section˜5.
Our novel decoding paradigm combines SAEs’ ability to disentangle superposed subspaces with the interpretability of conventional probes, overcoming DLA’s vocabulary dependence and feature-naming of SAE-based analysis. Although illustrated within a controlled testbed for input-competition tasks, our approach readily extends to other experimental settings, such as the question-answering scenario described in Section˜5.3. Appendix˜H discusses further applications, including bias and toxicity detection. Our probe reveals non-trivial insights into LLM representations, from GPT-2’s context-related richness to the richer embeddings for linguistic analogies. Our VSA-based probing paradigm is computationally efficient, with a lightweight probe that inspects a wide range of the model’s residual stream at minimal memory cost (see also Appendix˜Q). It applies to any autoregressive transformer, and its implementation works with any Hugging Face language models.
Additionally, the absence of theoretical limits in VSAs regarding the types of data, with the potentiality of hyperdimensional algebra, enables decoding multimodal latent features from LLM vector space. For example, the proof of concept in Appendix˜P shows VSA-based probing for a MNIST-based mathematical analogy (LeCun et al., 2010) and a VSA encoding for a multimodal input from the COCO dataset (Lin et al., 2014).
@@
The main limitation of our work (Appendix˜A) is its reliance on a predefined set of concepts. While we applied several strategies to mitigate confounding effects in probe learning, such as testing on syntactically different inputs, we could not measure their effectiveness. Section˜F.1 presents two control tests (Hewitt and Liang, 2019) to evaluate confounding effects on decoding.
@@ Appendix A Limitations
Table˜1, Table˜8 and Table˜9 report on the actual concepts identified by our probing method. The label Key values denotes instances where the probe retrieves a key of an analogy pair with a concept linked to it in a different domain (see Australia in Section˜4.1). This outcome can be viewed as an artifact of our probe, revealing the confounding influence of memorized key-value associations. Nevertheless, such cases constitute only a small fraction, 2% of the 114,099 textual inputs processed across all models, covering Key | Key Values, Example Value | Key Values, and Key Values | Target shown in Table˜1. To further investigate potential confounding effects from probe learning, we introduce two control tests, as proposed in [Hewitt and Liang, 2019]. Using randomly-permuted input embeddings ($e_{s}$) as a null model [Gotelli et al., 2012], and applying the unbinding operation on VSA encodings ($y_{s}$) with concept pairs unrelated to inputs, respectively, permuted and unrelated baseline in Appendix˜F.
In Section˜5.1, we introduce the value of 0.5 in the precision metric for LLM’s next-token prediction due to tokenization. While this approach is suitable for our needs, it may underestimate the model’s performance, resulting in lower precision@k scores in next-token prediction. However, in our experiments, only 8.6% of all instances fail in this scenario for the next token@1, considering the averaged ratio across models (Llama 4: 0.7%, Pythia: 16.2%; see also Section˜L.1 and Section˜L.2). On the other hand, models’ tokenizers could introduce variability in language models by themselves.
@@ D.1 Training details
The neural VSA encoder $\mathcal{M}$ was trained for 421 epochs on average via PyTorch Lighting,888lightning.ai/docs/pytorch/stable using early stopping (patient set at 100 epochs) and a batch size of 32. The optimal learning rate was automatically determined using the learning rate finder provided by the aforementioned library, and was approximately set to $3e^{-5}$ on average. We use AdamW as the optimizer (weight decay of $1e^{-4}$), applying a learning rate schedule based on Cosine Annealing with Warm Restarts, starting from the 100th epoch and doubling the restart period thereafter. To adapt the batch size after LR restarts, we employed a Gradient Accumulation Scheduler: the effective batch size was doubled at the 110th epoch, quadrupled at the 310th, and increased eightfold at the 410th epoch. During training, the model’s outputs are dynamically binarized using the sigmoid function to ensure compatibility with the loss function (Section˜4.4). This approach demonstrated better empirical performance than linear min-max normalization.
@@ Appendix F Experimental results
Table 5: Experimental results on the LLM’s next-token prediction, along with our probing method for retrieving the target concept from model’s latent representations. The model are ordered based on precision@1 for next-token prediction, and standard deviation is reported for each. To control for randomness, we also introduce two control tests using Llama 3.1-8B: a comparison against a null model with randomly-permuted input embeddings ($e_{s}$, permuted baseline), and extraction of concept pairs unrelated to inputs ($y_{s}$, unrelated baseline).
MODEL Next Token Prediction VSA-based Probing
Precison@1 Precison@5 Precison@1 Precison@5
Permuted baseline - - 0.080 $\pm$ 0.27 0.103 $\pm$ 0.30
Unrelated baseline - - 0.099 $\pm$ 0.30 0.105 $\pm$ 0.31
Llama 4 Scout, 17B-16E 0.077 $\pm$ 0.26 0.463 $\pm$ 0.48 0.866 $\pm$ 0.34 0.875 $\pm$ 0.33
GPT-2, medium 0.227 $\pm$ 0.39 0.471 $\pm$ 0.46 0.692 $\pm$ 0.46 0.702 $\pm$ 0.46
Pythia-1.4b 0.288 $\pm$ 0.41 0.541 $\pm$ 0.44 0.778 $\pm$ 0.42 0.790 $\pm$ 0.41
Llama 3.1-8B 0.309 $\pm$ 0.44 0.490 $\pm$ 0.47 0.891 $\pm$ 0.31 0.908 $\pm$ 0.29
Phi 4 0.478 $\pm$ 0.48 0.683 $\pm$ 0.43 0.887 $\pm$ 0.32 0.904 $\pm$ 0.30
OLMo-2 0.482 $\pm$ 0.48 0.656 $\pm$ 0.44 0.879 $\pm$ 0.33 0.890 $\pm$ 0.31
AVERAGE 0.310 $\pm$ 0.41 0.551 $\pm$ 0.45 0.832 $\pm$ 0.35 0.845 $\pm$ 0.33
@@ F.1 Validation strategy
To assess the effectiveness of our probe, we conduct two control tests (see Table˜5) as proposed in “Designing and interpreting probes with control tasks” by Hewitt and Liang [2019]:
@@ F.2 Distribution of instances with no concepts extracted
We examine probe performance across different LLM input types, defining success and failure by the presence or absence of concepts extracted by VSA probing. Table˜6 displays the distribution of instances with no concept extracted grouped by input categories. While we observe model-wise variability, this preliminary analysis shows a common pattern in representation blankness.
@@ Appendix G Experimental comparison
Our neural VSA encoder (Section˜4.4) does qualify as a supervised probe, as it is trained to map LLM internal representations (i.e., residual stream activations) into interpretable, human-understandable features (i.e., VSA encodings). Supervised probes are typically designed for specific experimental goals or target features, ranging from syntactic structure, as in “A Polar Coordinate System Represents Syntax in Large Language Models” [Diego Simon et al., 2024]; to real-world knowledge, as in “Language Models Represent Space and Time” [Gurnee and Tegmark, 2023]; and to abstract semantics, as in “The Geometry of Truth” [Marks and Tegmark, 2023]. Our probe is specifically designed around VSA principles, so direct comparisons with non-VSA probes would require fundamentally different approaches not grounded in VSAs.
@@ G.1 DLA-based experimental results
We adopt simple, fuzzy token-to-concept matching approach with our concept set (e.g., pes $\mapsto$ peso), and consider projected next-token predictions (Section˜G.3) from the model’s middle to last layers of the last token, as VSA probing. DLA produces no concepts in nearly 30% of analogies on average (see NONE in Table˜7; +17% compared to VSA, Table˜1), while yielding the target with its key in 26% of the cases (-50%). In instances without concepts from DLA, our VSA-based probe extracts, on average, the key-target pair in 57% of all analogies (Table˜8), while returning none for 28%. For instance, for the analogy king is to queen as son is to $\mapsto$ daughter, using OLMo-2, our probe extracts the key-target concepts (son and daughter), while DLA produces no concepts. The model predicts the next token prediction as ? with a softmax score of 0.06, followed by father (0.05); the target word has a rank of 37. Focusing on next-token representations, and thus capturing surface-level features, DLA exhibits inferior probing capabilities compared to ours, which compromise subsequent interpretability analyses of LLM embeddings. On the other hand, we observe substantial variance within this subset during VSA probing. Across models (Table˜8), our probe fails to retrieve any concepts in 43% of cases for Llama 4, compared to only 14% for Llama 3.1. GPT-2 confirms greater representativeness for the in-context example. There is also variation across analogy categories in this subset (Table˜9): for OLMo-2, linguistic analogies show the highest retrieval rates for Context $\mid$ Target (7.4% and 4.4%), whereas mathematical analogies shows nearly no concept retrieval (91%), confirming common blank representations. Section˜G.2 shows that, in cases where VSA fails, also DLA frequently yields no concepts rather than other relevant concepts.
@@
GPT-2 Pythia Llama 3.1 Phi-4 OLMo-2 Llama 4, Scout AVERAGE
DLA failures (docs, %) 33.9 32.8 47.4 33.1 14.6 15.4 29.5 $\pm$ 11.4
Concepts extracted by VSA (docs, %)
Key $\mid$ Target 53.5 56.5 76.6 70.5 44.5 42.6 57.4 $\pm$ 12.5
NONE 26.8 24.3 13.7 18.6 41.0 43.2 27.9 $\pm$ 10.9
Example 6.8 3.0 2.1 2.1 3.4 2.0 3.2 $\pm$ 1.7
Key 5.2 7.3 0.7 0.9 1.6 3.0 3.1 $\pm$ 2.4
Out-of-context 2.0 1.8 1.1 1.9 3.8 4.5 2.5 $\pm$ 1.2
Key $\mid$ Pair Values 1.8 2.1 2.5 3.2 2.4 1.8 2.3 $\pm$ 0.5
Context $\mid$ Target 0.7 1.2 1.1 0.7 1.9 1.4 1.2 $\pm$ 0.4
Target 0.1 0.1 0.2 0.0 0.1 0.0 0.1 $\pm$ 0.1
@@ H.1 Generalization of input representation
Scalability challenges depend largely on the nature of the input features. For tasks such as toxicity detection, expert-labeled data or specialized feature extraction pipelines may be required. For example, mapping the phrase “You are a pathetic excuse for a human just like the rest of your kind” to a conceptual form such as $(\phi_{\textrm{attack}}\odot\phi_{\textrm{insult}})+(\phi_{\textrm{attak}}\odot\phi_{\textrm{identity}})$ requires human expertise. Once features are extracted, however, constructing VSA encodings is automatic, efficient, and scalable. VSA probing can then uncover encoded concepts in the LLM vector space, for instance:

### v2

@@ 6 From input-completion tasks to text generation
This dataset evaluates extractive QA, where each answer is a text span within a given context, through questions generated by crowdworkers over Wikipedia articles. This aligns with our concept-focused probing, as questions and answers target contextual concepts, allowing us to benchmark feature extraction against features derived from both. We extract question- and answer-related concepts grounded in lexical semantics by identifying nouns, verbs, and adjectives using WordNet (Miller,, 1995) and DBpedia (Lehmann et al.,, 2015). Using bundling operations, we construct 693,886 training examples $\mathcal{Q}$ by progressively pairing SQuAD questions with their associated lexical features:
$\text{``What was the {name}''}\mapsto\phi_{\textrm{name}}$ (5)
$\text{``What was the {name} of the {ship}''}\mapsto(\phi_{\textrm{name}}+\phi_{\textrm{ship}})$
@@
Our trained encoder (Section˜4.4) achieves a test-set cosine similarity of 0.44, and a binary accuracy of 0.70 for mapping embeddings of Llama 3.1 into these VSA encodings. For our probing experiments, we consider 10,000 sampled items $\bar{\mathcal{Q}}$ from SQuAD, each formatted as a question preceded by its contextual text. See Appendix˜I for details. We then apply our probe to the model’s state both before and after text generation, comparing VSA encodings with the codebook $\Phi$ using cosine similarity with a concept detection threshold of 0.1, as unbinding is not required. Our probe extracts an average of three concepts before both the first and last token generation. The LLM’s QA-related performance achieved an average token-based F1 score of 0.69 (95% CI: 0.68–0.70), an exact match of 0.52 (95% CI: 0.51–0.53), and 68% of outputs (95% CI: 67–69) contained the target answers.
Observing concept drift in generation (RQ6).
We evaluate semantic-based concept relevance by computing cosine similarity between concept embeddings and question-answer features. The average similarity of extracted concepts related to the question decreases after text generation: by 4.8% for the entire sample and by 8.0% in the LLM error subset (32% of the sample), while no significant differences are observed prior to generation (Figure˜5; left). For answer-related concepts, overall no change is observed before and after text generation, but the LLM error subset shows a slight increase (+3.2–3.5%; Figure˜5, right). This suggests that LLM failures may stem from losing focus on the question rather than from a lack of answer-related knowledge. This hypothesis is supported by a weak positive Spearman correlation $(0.2$ with a p-value of $1e^{-99}$; Appendix˜K) between LLM’s F1 score and the proportion of question-related concepts extracted after text generation. For example, for the SQuAD query “What do laboratories try to produce hydrogen from?” (target answer: “solar energy and water”), the model erroneously outputs “water and heat” (F1 = 0.57). Before the model’s text generation, our proposed approach extracts the concepts $\phi_{\textrm{try}}$, $\phi_{\textrm{produce}}$, $\phi_{\textrm{hydrogen}}$ (question) and $\phi_{\textrm{solar}}$, $\phi_{\textrm{water}}$ (answer); after generation, the question-related concept set reduced to $\phi_{\textrm{produce}}$ and the answer set gained the concept $\phi_{\textrm{energy}}$. During text generation, the model lost the concept of hydrogen while refining concepts related to answering. This leads to a response that no longer focuses on the question’s subject, but reflect the more general concepts still active in the model’s internal state.
@@ 7 Conclusions
This work provides empirical evidence that the latent features of neural embeddings can be faithfully captured using VSA-based representations (RQ7; Section˜4.4). Building on this, VSA probing enables a unified, concept-oriented analysis of features aligned with both inputs and outputs (RQ8; Section˜5.2, Section˜6); integrating the complementary viewpoints of prior approaches for probing.
Our method combines the top-down interpretability of supervised probes (Section˜4.2), SAE’s sparsity-driven proxy space (Section˜4.4), and output-oriented investigation (Section˜4.5) of logit-based methods. Additionally, our ingestion algorithm (Section˜4.3) bypasses layer selection of conventional layer-wise approaches. It enables joint input–output feature extraction that uncovers non-trivial insights into neural embeddings. Section˜5.2 shows differences in conceptual richness (RQ1–RQ3), including OLMo-2’s richer structure compared to GPT-2 (RQ5), and the common sparsity in mathematical analogies (RQ4). Section˜6 reveals concept-related patterns during text generation (RQ6); concept extraction before and after generation suggests that LLM failures may stem from losing focus on question-related concepts rather than answer-related ones.
Our methodology applies to any autoregressive model, is compatible with all Hugging Face models, and introduces a layer-agnostic, lightweight probe (Appendix˜Q). Moreover, combining VSA generality with hypervector algebra provides a promising way to investigate multimodal features in neural embeddings, as displayed in the proof of concept in Appendix˜P.
@@
The primary limitations of our work are its dependence on unidirectional transformation that prevents causal validation and on a set of predefined features (Appendix˜A). While we implemented multiple strategies to reduce confounding effects in probe-induced learning, such as evaluating on syntactically diverse inputs ($\bar{\mathcal{S}}$, and $\bar{\mathcal{Q}}$), we were unable to directly measure their effectiveness. Section˜F.1 shows two validation tests to assess confounding effects on decoding.
@@ Appendix A Limitations
Table˜1, Table˜2 and Table˜9 report on the actual concepts identified by our probing method. The label Key values denotes instances where the probe retrieves a key of an analogy pair with a concept linked to it in a different domain (see Australia in Section˜4.1). This outcome can be viewed as an artifact of our probe, revealing the confounding influence of memorized key-value associations. Nevertheless, such cases constitute only a small fraction, 2% of the 114,099 textual inputs processed across all models, covering Key | Key Values, Example Value | Key Values, and Key Values | Target shown in Table˜1. To further investigate potential confounding effects from probe learning, we introduce two control tests, as proposed in (Hewitt and Liang,, 2019). Using randomly-permuted input embeddings ($e_{s}$) as a null model (Gotelli et al.,, 2012), and applying the unbinding operation on VSA encodings ($y_{s}$) with concept pairs unrelated to inputs, respectively, permuted and unrelated baseline in Appendix˜F.
@@ D.1 Training details
The neural VSA encoder $\mathcal{M}$ was trained for 421 epochs on average via PyTorch Lighting,777lightning.ai/docs/pytorch/stable using early stopping (patient set at 100 epochs) and a batch size of 32. The optimal learning rate was automatically determined using the learning rate finder provided by the aforementioned library, and was approximately set to $3e^{-5}$ on average. We use AdamW as the optimizer (weight decay of $1e^{-4}$), applying a learning rate schedule based on Cosine Annealing with Warm Restarts, starting from the 100th epoch and doubling the restart period thereafter. To adapt the batch size after LR restarts, we employed a Gradient Accumulation Scheduler: the effective batch size was doubled at the 110th epoch, quadrupled at the 310th, and increased eightfold at the 410th epoch. During training, the model’s outputs are dynamically binarized using the sigmoid function to ensure compatibility with the loss function (Section˜4.4). This approach demonstrated better empirical performance than linear min-max normalization.
@@ Appendix F Experimental results
Table 6: Experimental results on the LLM’s analogy-style completion tasks, along with our probing method for retrieving the target concept from model’s internal state. The model are ordered based on precision@1 for next-token prediction. Statistical variability is reported using 95% confidence intervals, which more appropriately capture variability for metrics bounded within the [0,1] range. To control for randomness, we also introduce two control tests using Llama 3.1-8B: a comparison against a null model with randomly-permuted input embeddings ($e_{s}$, permuted baseline), and extraction of concept pairs unrelated to inputs ($y_{s}$, unrelated baseline).
MODEL LLM’s Next Token DLA-based Probing VSA-based Probing
Precision@1 Precision@5 Precision@1 Precision@1 Precision@5
Permuted baseline - - - 0.080 (079-082) 0.103 (101-104)
Unrelated baseline - - - 0.099 (097-101) 0.105 (103-107)
Llama 4 Scout, 17B-16E 0.081 (079-082) 0.501 (498-504) 0.773 (771-775) 0.866 (864-868) 0.875 (873-877)
GPT-2, medium 0.267 (265-270) 0.548 (545-551) 0.300 (298-303) 0.692 (689-695) 0.702 (699-704)
Pythia-1.4b 0.369 (366-372) 0.654 (651-657) 0.413 (410-416) 0.778 (776-781) 0.790 (788-792)
Llama 3.1-8B 0.352 (349-354) 0.546 (543-549) 0.467 (464-470) 0.891 (889-893) 0.908 (907-910)
Phi 4 0.519 (516-522) 0.753 (750-755) 0.585 (582-588) 0.887 (886-889) 0.904 (902-905)
OLMo-2 0.529 (526-532) 0.721 (719-724) 0.714 (712-717) 0.879 (877-881) 0.892 (890-894)
AVERAGE 0.352 0.621 0.542 0.832 0.845
@@ F.1 Validation strategy
To assess the effectiveness of our probe, we conduct two control tests (see Table˜6) as proposed in “Designing and interpreting probes with control tasks” by Hewitt and Liang, (2019):
@@ F.2 Distribution of instances with no concepts extracted
We examine probe performance across different LLM input types, defining success and failure by the presence or absence of concepts extracted by VSA probing. Table˜7 displays the distribution of instances with no concept extracted grouped by input categories. While we observe model-wise variability, this preliminary analysis shows a common pattern in representation blankness.
@@ Appendix G Experimental comparison
Our neural VSA encoder (Section˜4.4) does qualify as a supervised probe, as it is trained to map LLM internal representations (i.e., residual stream activations) into interpretable, human-understandable features (i.e., VSA encodings). Supervised probes are typically designed for specific experimental goals or target features, ranging from syntactic structure, as in “A Polar Coordinate System Represents Syntax in Large Language Models” (Diego Simon et al.,, 2024); to real-world knowledge, as in “Language Models Represent Space and Time” (Gurnee and Tegmark,, 2023); and to abstract semantics, as in “The Geometry of Truth” (Marks and Tegmark,, 2023). Our probe is specifically designed around VSA principles, so direct comparisons with non-VSA probes would require fundamentally different approaches not grounded in VSAs.
@@ G.1 DLA-based experimental results
We adopt simple, fuzzy token-to-concept matching approach with our concept set (e.g., pes $\mapsto$ peso), and consider projected next-token predictions (Section˜G.3) from the model’s middle to last layers of the last token, as VSA probing. DLA produces no concepts in nearly 30% of analogies on average (see NONE in Table˜8; +17% compared to VSA, Table˜1), while yielding the target with its key in 26% of the cases (-50%). In instances without concepts from DLA, our VSA-based probe extracts, on average, the key-target pair in 57% of all analogies (Table˜2), while returning none for 28%. For instance, for the analogy king is to queen as son is to $\mapsto$ daughter, using OLMo-2, our probe extracts the key-target concepts (son and daughter), while DLA produces no concepts. The model predicts the next token prediction as ? with a softmax score of 0.06, followed by father (0.05); the target word has a rank of 37. Focusing on next-token representations, and thus capturing surface-level features, DLA exhibits inferior probing capabilities compared to ours, which compromise subsequent interpretability analyses of LLM embeddings. On the other hand, we observe substantial variance within this subset during VSA probing. Across models (Table˜2), our probe fails to retrieve any concepts in 43% of cases for Llama 4, compared to only 14% for Llama 3.1. GPT-2 confirms greater representativeness for the in-context example. There is also variation across analogy categories in this subset (Table˜9): for OLMo-2, linguistic analogies show the highest retrieval rates for Context $\mid$ Target (7.4% and 4.4%), whereas mathematical analogies shows nearly no concept retrieval (91%), confirming common blank representations. Section˜G.2 shows that, in cases where VSA fails, also DLA frequently yields no concepts rather than other relevant concepts.
@@ H.1 Generalization of input representation
Scalability challenges depend largely on the nature of the input features. For tasks such as toxicity detection, expert-labeled data or specialized feature extraction pipelines may be required. For example, mapping the phrase “You are a pathetic excuse for a human just like the rest of your kind” to a conceptual form such as $(\phi_{\textrm{attack}}\odot\phi_{\textrm{insult}})+(\phi_{\textrm{attack}}\odot\phi_{\textrm{identity}})$ requires human expertise. Once features are extracted, however, constructing VSA encodings is automatic, efficient, and scalable. VSA probing can then uncover encoded concepts in the LLM vector space, for instance: